OpenStudy (elleblythe):

Factor completely x^4+10x^2+49

OpenStudy (cwrw238):

r u sure you thats correct?

OpenStudy (elleblythe):

@cwrw238 yes, there's no typo

OpenStudy (rsadhvika):

its not x^4+14x^2+49 ?

OpenStudy (elleblythe):

@rsadhvika nope that's the given polynomial

OpenStudy (rsadhvika):

okay thank you :) put x^2 = t you will see a quadratic in t use quadratic formula and solve t

OpenStudy (anonymous):

unless you wanna go into imaginary form...x^4+10x^2+49=(x^2-2x+7)(x^2+2x+7)

OpenStudy (rsadhvika):

wolfram says the same, point is how to get to that

OpenStudy (anonymous):

imaginary